Evaluate the following integral using the partial fraction method
Jee - А) Зах — 1 dx х (х2 — 4).
Answer #1

B 32-1 0-1 B= 32-1 7 32-1 3K-1 A x (424) x (2-2) (472) A (x-2)(2+2) + 8 %(x+2) + C 2(3-2) x (x-2)(x+2) u (x-2)(x+2) 32-) = A